In an experiment of rolling a balanced die twice, let X be the maximum of the two numbers obtained.

1) What are the possible values of X?

2) Write down the sample space

3) Wat is the probability function for X?

4) Determine and sketch the probability mass function of X.

5) Find the expected value of X.

6) Find the variance of X.



1
Expert's answer
2022-05-24T17:52:06-0400

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 1 & 2 & 3 & 4 & 5 & 6 \\
 \hdashline
 2 & 2 & 2 & 3 & 4 & 5 & 6 \\
 \hdashline
 3 & 3 & 3 & 3 & 4 & 5 & 6 \\
 \hdashline
 4 & 4 & 4 & 4 & 4 & 5 & 6 \\
 \hdashline
 5 & 5 & 5 & 5 & 5 & 5 & 6 \\
 \hdashline
 6 & 6 & 6 & 6 & 6 & 6 & 6 \\
 \hdashline
\end{array}"

1) The possible values of "X" are "1, 2, 3, 4, 5," and "6."


2)

"S=\{1,2,3,4,5,6\}"

3)


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 p(x) & 1/36 & 3/36 & 5/36 & 7/36 & 9/36 & 11/36 \\

\end{array}"

4)


"f(x) = \begin{cases}
 1/36 &\ x=1 \\
 3/36&\ x=2 \\
 5/36&\ x=3 \\
 7/36&\ x=4 \\
 9/36&\ x=5 \\
 11/36&\ x=6 \\
\end{cases}"



5)


"\mu=E(X)=\dfrac{1}{36}(1)+\dfrac{3}{36}(2)+\dfrac{5}{36}(3)"

"+\dfrac{7}{36}(4)+\dfrac{9}{36}(5)+\dfrac{11}{36}(6)=\dfrac{161}{36}"

6)


"E(X^2)=\dfrac{1}{36}(1)^2+\dfrac{3}{36}(2)^2+\dfrac{5}{36}(3)^2"


"+\dfrac{7}{36}(4)^2+\dfrac{9}{36}(5)^2+\dfrac{11}{36}(6)^2=\dfrac{791}{36}"


"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=\dfrac{791}{36}-(\dfrac{161}{36})^2=\dfrac{2555}{1296}"